OpenStudy (anonymous):

What is the value of x in the equation −x = 3 − 4x + 6? −9 −3 3 9

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

um -9

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

its 9

OpenStudy (anonymous):

Let me draw it for you

OpenStudy (anonymous):

\[-x=3-4x+6\]

OpenStudy (anonymous):

You have to add 4x to each side

OpenStudy (anonymous):

um im sorry i dont understand

OpenStudy (anonymous):

\[3x=3+6\]

OpenStudy (anonymous):

There is an x on both sides of the equal sign. Correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So you would a 4x to -x

OpenStudy (anonymous):

and what would that be?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Exactly 3x

OpenStudy (anonymous):

yay thank you

OpenStudy (anonymous):

But that isn't the answer

OpenStudy (anonymous):

oh :/

OpenStudy (anonymous):

Add -4x to 4x

OpenStudy (anonymous):

That would make it 0.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[3x=3+6\]

OpenStudy (anonymous):

So you'll have that. 3+6 is?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Precisely.

OpenStudy (anonymous):

\[3x=9\]

OpenStudy (anonymous):

Then divide and that's your answer

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Good. Now I know you choose that answer before but I wanted to make sure that You new why that was the answer.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

Can I have a medal and a fan?

OpenStudy (anonymous):

If you need help on anything else, tag me